# SceneDreamer360: Text-Driven 3D-Consistent Scene Generation with Panoramic Gaussian Splatting

### v1

@@ Introduction
The scarcity of annotated 3D text-point cloud datasets presents significant challenges in training 3D point cloud models directly from user queries, especially for complex 3D scenes. This challenge arises primarily because the cost-intensive nature of acquiring and annotating 3D data, coupled with the high precision required for modeling intricate 3D scenes. Consequently, existing methods often extend high-quality 2D generative models by leveraging 2D object priors to transition into 3D neural radiance fields. For instance, CLIP-NeRF (Wang et al. 2022) employs the CLIP (Radford et al. 2021) model to control the shape and appearance of 3D objects via textual prompts or images. Dream Fields (Jain et al. 2022) integrates neural rendering with image and text representations to generate diverse 3D objects from natural language descriptions. Similarly, Dreamfusion (Poole et al. 2022a) optimizes NeRF without requiring 3D training data, relying solely on a pre-trained 2D text-to-image diffusion model for text-to-3D synthesis.
@@
Figure 1: We introduce SceneDreamer360, a text-based 3D scene generation framework designed to create realistic 3D scenes with high consistency across different viewpoints.
@@
Despite significant advancements in NeRF-based point cloud generation methods, generating consistent point clouds with fine-grained details remains a challenge. The Text2Room (Höllein et al. 2023) method often generates point clouds that are discontinuous, lack fine details, and require extended rendering times. Similarly, LucidDreamer (Chung et al. 2023) generates panoramic point clouds with numerous inconsistencies, resulting in incomplete panoramas. PanFusion (Zhang et al. 2024) also faces difficulties in generating panoramic images from complex long texts, leading to blurred areas and object deformations. Additionally, PanFusion’s panoramic images have a relatively low resolution ($512\times 1024$), leading to fewer spatial points corresponding to pixel points, which causes the rendered point clouds to appear blurred and fall short of ideal quality. To address these challenges, we propose introducing 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) for multi-scene generation, with a focus on producing more finely detailed complex scene point clouds. 3DGS utilizes Gaussian distribution for scene modeling, enabling finer representation of complex scene point clouds and enhancing reconstruction accuracy. Furthermore, with an optimized rendering algorithm, 3DGS significantly reduces rendering times while maintaining high quality, resulting in more efficient generation outcomes.
@@
In this paper, we propose a novel SceneDreamer360 for text-driven 3D-consistent scene generation using panoramic Gaussian splatting. As shown in Fig. 1, our approach uses a two-stage process: panoramic image enhancement for 3D consistency priors, followed by 3DGS to reconstruct high-quality, text-aligned point clouds. For panorama generation and enhancement, existing methods enhance image resolution through iterative progressive scene rendering, often leading to consistency issues between successive renderings. To address this, we fine-tune the PanFusion (Zhang et al. 2024) model by incorporating an MLP (Murtagh 1991) and LoRA (Hu et al. 2021) layer at the final stage, training it using the Habitat Matterport Dataset (Ramakrishnan et al. 2021). This approach yields a model checkpoint tailored to specific needs, enhancing the quality of panoramic image generation. We further enrich and upscale the panoramas using CN (Zhang, Rao, and Agrawala 2023) and RealESRGAN (Wang et al. 2021), achieving resolutions up to 6K. This enhancement improves point cloud rendering, preserving high-resolution details for more aesthetically pleasing and realistic results. During point cloud fusion, we designe a novel point cloud initialization method to enhance consistency and reduce rendering time. Additionally, leveraging the effective representation of 3DGS enables the generation of complete and detailed point cloud images. Our contributions can be summarised as follows:
@@
Related Work
@@
3D Scene Generation. The field of 3D scene generation has evolved significantly, drawing inspiration from various breakthroughs in image generation techniques. Early approaches leveraged Generative Adversarial Networks (GANs) (Goodfellow et al. 2014) , attempting to create multi-view consistent images (Chan et al. 2021; Niemeyer and Geiger 2021) or directly generate 3D representations like voxels (Nguyen-Phuoc et al. 2019, 2020) and point clouds (Achlioptas et al. 2018; Shu, Park, and Kwon 2019). Later, the advent of diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) and their success in image generation (Ramesh et al. 2021; Rombach et al. 2022) sparked a new wave of research in 3D scene generation. Researchers began applying diffusion models to various 3D representations, including voxels (Zhou, Du, and Wu 2021), point clouds (Vahdat et al. 2022), and implicit neural networks (Poole et al. 2022b). While these approaches showed promise, they often focused on simple, object-centric examples due to their inherent nature.
@@
To address more complex scenarios, some generative diffusion models employed meshes as proxies, diffusing in UV space. This approach enabled the creation of large portrait scenes through continuous mesh building (Fridman et al. 2024) and the generation of indoor scenes (Cohen-Bar et al. 2023; Lei, Tang, and Jia 2023) and more realistic objects (Qian et al. 2023). Currently, an increasing number of studies (Chung et al. 2023; Zhou et al. 2024b; Ma, Zhan, and Jin 2024; Li et al. 2024) have incorporated 3DGS (Kerbl et al. 2023) as a 3D representation to complement diffusion models in generating more complex 3D scenes. However, due to diffusion models generating one viewpoint at a time before merging and elevating to 3D space, the resulting 3D scenes often lack realism and consistency.
@@
In contrast, to address the issues of quality and spatial consistency, our method leverages the promising 3D Gaussian Splatting technique and implements a three-stage enhancement process for the generated panoramic images. Concurrent work HoloDreamer (Zhou et al. 2024a) has explored similar concepts. However, our approach is distinguished by an innovative design in the panoramic image enhancement stage and a novel point cloud fusion algorithm.
@@
3D Scene Representation. The field of 3D scene representation has evolved significantly, encompassing a variety of approaches with distinct trade-offs and applications. Traditional explicit methods such as point clouds, meshes, and voxels have long been fundamental to 3D modeling, offering intuitive control and fast rendering through rasterization. However, these methods often require a large number of elements to achieve high resolution, which can be computationally expensive. To address this limitation, more complex primitives are developed, including cuboids (Tulsiani et al. 2017) and polynomial surfaces (Yavartanoo et al. 2021), offering more efficient expression of geometry but still struggling with realistic color representation.
@@
Recent advancements have led to the emergence of implicit neural representations, such as Signed Distance Functions (SDF) (Takikawa et al. 2021) and Neural Radiance Fields (NeRF) (Mildenhall et al. 2021). Among these advancements, 3D Gaussian Splatting(3DGS) (Kerbl et al. 2023) has emerged as a particularly promising approach. It represents complete and unbounded 3D scenes by effectively ’splatting’ Gaussians, utilizing spherical harmonics and opacity for strong representational capabilities. 3DGS balances quality and efficiency by supporting alpha-blending and differentiable rasterization, enabling fast 3D scene optimization. It refines point clouds details through a split-and-clone mechanism, making it ideal for complex scene generation where scene bounds are uncertain.
@@
Panorama Generation Panoramas offer a comprehensive, unobstructed view that captures extensive scene areas. Recent advancements in panorama generation have significantly improved the quality and consistency of 360-degree images. While early attempts like PanoGen (Li and Bansal 2024) and MultiDiffusion (Bar-Tal et al. 2023) used pre-trained diffusion models to create long images from text prompts, these methods often resulted in inconsistencies and lacked true panoramic properties. To address these limitations, researchers have developed more sophisticated approaches. MVDiffusion (Tang et al. 2023) introduced a correspondence-aware attention mechanism to generate consistent multi-view images simultaneously. StitchDiffusion (Wang et al. 2024) improved upon it by using LoRA (Hu et al. 2021) fine-tuning to generate complete 360-degree panoramas with enhanced continuity between image edges. PanFusion (Zhang et al. 2024)proposed a novel dual-branch diffusion model to generate a 360-degree image from a text prompt. In this paper, we incorporated MLP and LoRA into the PanFusion model and performed fine-tuning to enhance the consistency of the generated panoramic images.
Method
This method introduces a novel pipeline for generating high-resolution 360-degree panoramic point cloud scenes from arbitrary spatial text, consisting of two main stages: Panoramic Image Generation and Processing, and Point Cloud Construction with Gaussian Splatting. The overall pipeline of SceneDreamer360 is shown in Fig. 2 .
Panoramic Image Generation and Processing
@@
Panoramic Image Generation. Our method utilized the PanFusion model (Zhang et al. 2024) to generate panoramas. However, because the PanFusion model was originally trained on simple short text prompts, it required appropriate adjustments and fine-tuning to handle complex long texts. Therefore, we added a Multi-Layer Perceptron (MLP) (Murtagh 1991) layer and a Low-Rank Adaptation (LoRA) (Hu et al. 2021) layer to the original PanFusion model, fine-tuning it using the reannotated Habitat Matterport Dataset (Ramakrishnan et al. 2021). The MLP layer captures higher-order features in complex long texts, while the LoRA layer enhances the model’s representation ability through low-rank matrix decomposition. For model fine-tuning, we first reannotated the dataset using the generated complex long texts, pairing them with corresponding panoramas to form new training samples. Next, we froze the pre-trained parts of the PanFusion model and trained only the newly added MLP and LoRA layers. This ensures that the model effectively utilizes information from complex long texts to generate high-quality panoramas. These panoramas encapsulate the details and scenes described in the input text, serving as foundational data for subsequent steps.
@@
Panoramic Image Optimization. As the pre-trained PanFusion model can only generate panoramas with a resolution of $512\times 1024$, directly using these panoramas for multi-view extraction and point cloud reconstruction would result in sparse point clouds with blurry 3D Gaussian rendering (Kerbl et al. 2023). To meet the requirements of point cloud reconstruction, it is necessary to further enhance the resolution and quality of the panoramas. The specific steps are as follows: We adopted a three-step super-resolution reconstruction method involving ControlNet-Tile (Zhang, Rao, and Agrawala 2023), RealESRGAN (Wang et al. 2021), and a second application of ControlNet-Tile to maximize the enhancement of image details and resolution.

### v2

@@ I Introduction
The scarcity of annotated 3D text-point cloud datasets poses substantial challenges to training 3D point cloud models directly from user queries, particularly for complex 3D scenes. This difficulty primarily stems from the high costs associated with acquiring and annotating 3D data, in addition to the precision required for modeling detailed 3D environments. To address these limitations, existing approaches frequently extend advanced 2D generative models by leveraging 2D object priors as a bridge to 3D neural radiance fields. For instance, CLIP-NeRF [1] uses the CLIP model [2] to control the shape and appearance of 3D objects based on textual prompts or images. Similarly, Dream Fields [3] integrates neural rendering with image and text representations to produce diverse 3D objects from natural language descriptions. Dreamfusion [4] further advances this approach by optimizing NeRF without the need for 3D training data, instead relying solely on a pre-trained 2D text-to-image diffusion model for text-to-3D synthesis.
@@
Figure 1: We introduce SceneDreamer360, a text-based 3D scene generation framework designed to create realistic 3D scenes with high consistency across different viewpoints. SceneDreamer360 consists of two stages. In the first stage, the Panorama Generation module creates enhanced panoramic images to provide a complete and consistent spatial prior for the 3D scene. In the second stage, 3D Gaussian Splatting is used to reconstruct multi-view spatial images, resulting in a complete and spatially consistent point cloud.
@@
Despite significant advancements in NeRF-based point cloud generation methods, achieving consistent, fine-grained point cloud details remains a challenge. For example, the Text2Room method [5] often produces point clouds that are discontinuous, lack detailed features, and require prolonged rendering times. Similarly, LucidDreamer [6] generates panoramic point clouds with numerous inconsistencies, resulting in incomplete and fragmented panoramas. PanFusion [7] also struggles with generating high-quality panoramic images from complex, extended text inputs, frequently yielding blurred areas and deformed objects. Moreover, PanFusion’s panoramic images are relatively low-resolution ($512\times 1024$), leading to a reduced spatial-to-pixel point ratio, which further contributes to blurred, suboptimal point cloud rendering. To address these challenges, we propose incorporating 3D Gaussian Splatting (3DGS) [8] for multi-scene generation, specifically aimed at producing more finely detailed and consistent complex scene point clouds. By employing a Gaussian distribution for scene modeling, 3DGS enables more precise representation of complex structures and improves reconstruction accuracy. Additionally, 3DGS’s optimized rendering algorithm significantly reduces rendering times while preserving high quality, resulting in faster and more efficient generation of detailed point clouds.
@@
In this paper, we introduce SceneDreamer360, a novel framework for text-driven 3D-consistent scene generation using panoramic Gaussian splatting (3DGS). Our method operates in two stages, as illustrated in Fig. 1: first, we enhance panoramic images to establish 3D consistency priors, and second, we apply 3DGS to reconstruct high-quality, text-aligned point clouds. This dual-stage approach ensures both accurate 3D structure and enhanced visual quality and realism in the generated scenes. During the panorama generation and enhancement stage, existing techniques frequently rely on iterative, progressive scene rendering, which can lead to consistency issues across successive renderings. To address this challenge, we enhance the PanFusion model [7] by integrating a multi-layer perceptron (MLP) [9] and a LoRA layer [10] into its final processing stage. These additions, combined with training on the Habitat Matterport Dataset [11], produce a model checkpoint fine-tuned to the specific demands of panoramic image generation. This enhancement step produces high-quality panoramas, forming a solid foundation for subsequent point cloud rendering. Further, we upscale and refine the generated panoramas using ControlNet (CN) [12] and RealESRGAN [13] techniques, achieving resolutions up to 6K. This upscaling process preserves fine details and enhances image fidelity, which is crucial for realistic point cloud rendering. High-resolution panoramas contribute to more detailed and visually appealing scenes, which are important for achieving photorealism in the final output. To optimize the point cloud rendering process, we introduce a novel point cloud initialization method that improves 3D consistency and reduces rendering time. This step ensures better alignment of the resulting point clouds with the panoramic views, creating a more cohesive and immersive 3D experience. Additionally, the robust 3D representation capabilities of 3DGS enable the creation of complete, high-quality point cloud images that are consistent with the input text prompts. Our contributions can be summarised as follows:
@@
The remainder of this paper is organized as follows: Section II reviews the background and related work on 3D scene generation, 3D scene representation, and panorama generation. Section III describes the architecture and functionality of the proposed Scenedreamer360 model in detail. Section IV presents the experimental results and visualizations, highlighting the effectiveness of our approach. Finally, Section V summarizes the study’s main findings and contributions.
II Related Work
II-A 3D Scene Generation
The field of 3D scene generation has evolved significantly, drawing inspiration from various breakthroughs in image generation techniques. Early approaches leveraged Generative Adversarial Networks (GANs) [14], attempting to create multi-view consistent images [15, 16] or directly generate 3D representations like voxels [17, 18] and point clouds [19, 20]. However, these methods were hindered by GAN’s inherent learning instability and the memory constraints of 3D representations, limiting the quality of generated scenes.
@@ II Related Work
The advent of diffusion models [21, 22, 23] and their success in image generation [24, 25] sparked a new wave of research in 3D scene generation. Researchers began applying diffusion models to various 3D representations, including voxels [26], point clouds [27], and implicit neural networks [28]. While these approaches showed promise, they often focused on simple, object-centric examples due to their inherent nature.
@@
To address more complex scenarios, some generative diffusion models employed meshes as proxies, diffusing in UV space. This approach enabled the creation of large portrait scenes through continuous mesh building [29] and the generation of indoor scenes [30, 31] and more realistic objects [32]. Currently, an increasing number of studies [6, 33, 34, 35] have incorporated 3DGS [8] as a 3D representation to complement diffusion models in generating more complex 3D scenes. For instance, LucidDreamer [6] pioneered the integration of 3D Gaussian Splatting into scene generation, demonstrating the efficacy of 3DGS. RealmDreamer [36] optimizes a 3D Gaussian Splatting representation to align with complex text prompts. However, due to diffusion models generating one viewpoint at a time before merging and elevating to 3D space, the resulting 3D scenes often lack realism and consistency. To address this issue, some recent methods such as DreamScene360 [33] and FastScene [34] have introduced panoramic images to generate more coherent 3D scenes. Nevertheless, these approaches still produce scenes lacking in detail and quality, primarily due to limitations in panoramic image generation models and view fusion algorithms.
@@
In contrast, to address the issues of quality and spatial consistency, our method leverages the promising 3D Gaussian Splatting technique and implements a three-stage enhancement process for the generated panoramic images. Concurrent work HoloDreamer [37] has explored similar concepts. However, our approach is distinguished by an innovative design in the panoramic image enhancement stage and a novel point cloud fusion algorithm. Our approach improves the overall quality and spatial coherence of the generated 3D scenes, advancing the state-of-the-art in text-driven 3D scene generation.
II-B 3D Scene Representation
@@
The field of 3D scene representation [38, 39] has evolved significantly, encompassing a variety of approaches with distinct trade-offs and applications. Traditional explicit methods such as point clouds, meshes, and voxels have long been fundamental to 3D modeling, offering intuitive control and fast rendering through rasterization. However, these methods often require a large number of elements to achieve high resolution, which can be computationally expensive. To address this limitation, more complex primitives are developed, including cuboids [40, 41, 42] and polynomial surfaces [43], offering more efficient expression of geometry but still struggling with realistic color representation.
Recent advancements have led to the emergence of implicit neural representations, such as Signed Distance Functions (SDF) [44] and Neural Radiance Fields (NeRF) [45]. These methods, particularly NeRF, have demonstrated the ability to represent complex 3D shapes and textures with rich details. However, they face challenges in terms of handling and optimization efficiency. To improve upon these limitations, subsequent research has focused on combining volume rendering with explicit representations, utilizing structures like sparse voxels [46, 47, 48, 49], featured point clouds [50], Multi-Level Hierarchies [51, 52], tensors [53], infinitesimal networks [54, 55], triplanes [56], and polygons [57].
@@
Among these advancements, 3D Gaussian Splatting(3DGS) [8] has emerged as a particularly promising approach. It represents complete and unbounded 3D scenes by effectively ’splatting’ Gaussians, utilizing spherical harmonics and opacity for strong representational capabilities. 3DGS balances quality and efficiency by supporting alpha-blending and differentiable rasterization, enabling fast 3D scene optimization. It refines point clouds details through a split-and-clone mechanism, making it ideal for complex scene generation where scene bounds are uncertain.
@@
II-C Panorama Generation
Panoramas offer a comprehensive, unobstructed view that captures extensive scene areas. Recent advancements in panorama generation have significantly improved the quality and consistency of 360-degree images. While early attempts like PanoGen [58] and MultiDiffusion [59] used pre-trained diffusion models to create long images from text prompts, these methods often resulted in inconsistencies and lacked true panoramic properties. To address these limitations, researchers have developed more sophisticated approaches. MVDiffusion [60] introduced a correspondence-aware attention mechanism to generate consistent multi-view images simultaneously. This method ensures that the different views maintain coherence and alignment, which is crucial for creating seamless panoramic images. However, MVDiffusion still faced challenges in achieving perfect continuity at the image edges. StitchDiffusion [61] improved upon it by using LoRA [10] fine-tuning to generate complete 360-degree panoramas with enhanced continuity between image edges. This approach leverages the Low-Rank Adaptation (LoRA) technique to fine-tune the model, ensuring that the transitions between different parts of the panorama are smooth and natural. PanFusion [7] proposed a novel dual-branch diffusion model to generate a 360-degree image from a text prompt. This model utilizes two separate branches to handle different aspects of the image generation process, allowing for more detailed and context-aware outputs. The dual-branch architecture helps in capturing both the global structure and local details of the scene, resulting in high-quality panoramic images.
In this paper, we incorporated MLP and LoRA into the PanFusion model and performed fine-tuning to enhance the consistency of the generated panoramic images. By integrating a Multi-Layer Perceptron (MLP) with LoRA in the PanFusion framework, we aim to further improve the coherence and quality of the panoramas. The MLP component helps in refining the details and ensuring that the generated images are not only consistent but also rich in detail. Furthermore, recent studies have explored the integration of machine learning techniques with traditional computer vision methods to further enhance panorama generation. For instance, hybrid models that combine deep learning with geometric alignment techniques have shown promise in reducing artifacts and improving the overall visual quality of the panoramas. These advancements highlight the ongoing efforts to push the boundaries of what is possible in panorama generation, making it an exciting and rapidly evolving field. Overall, the integration of advanced attention mechanisms, fine-tuning techniques, and hybrid models has significantly advanced the state of panorama generation, enabling the creation of high-quality, seamless 360-degree images from text prompts.
III Method
This method introduces a novel pipeline for generating high-resolution, 360-degree panoramic point cloud scenes from arbitrary spatial text inputs. The pipeline consists of two main stages: Panoramic Image Generation and Processing and Point Cloud Construction with Gaussian Splatting. As illustrated in Fig. 2, the SceneDreamer360 framework effectively combines these stages to produce visually coherent and spatially consistent 3D scenes, capturing fine details and maintaining high image fidelity.
In the first stage, an adapted PanFusion [7] model generates initial panoramic images, which are subsequently refined through a three-step super-resolution process, enhancing both image details and resolution. Next, we use the high-resolution panorama as input and apply the Equirectangular to Perspective (e2p) algorithm to obtain multiple scene images from different perspectives. In the second stage, point clouds are constructed using a 3D Gaussian Splatting [8] technique. This process involves initializing the point cloud, applying masking to remove duplicates for a more complete map, and training a 3D Gaussian splatting model along a new rendering trajectory to produce a detailed 3D spatial scene point cloud.
III-A Panoramic Image Generation and Processing
@@
III-A1 Complex Long-Text Generation
We used the GPT-4 model to generate a series of similar complex scene descriptions to increase the diversity and richness of the training data. The user first inputs an initial long text description of a complex scene. To generate more texts with similar semantics and content, we utilize GPT-4 to expand and transform it. These generated texts cover different details and perspectives, making the subsequent panorama generation more diverse and comprehensive.
III-A2 Finetune PanFusion Model
Given the input long text representation $T$ and its embedding $e_{T}$, the representation after the MLP layer is:
$\mathbf{h}=\sigma(\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\cdot\mathbf{e}_{T}+\mathbf{b}_{1})+\mathbf{b}_{2}),$ (1)
where $W_{1}$ and $W_{2}$ are the weight matrices of the MLP layer, $b_{1}$ and $b_{2}$ are the bias terms, and $\sigma$ is the activation function.
III-B Panoramic Image Generation
@@
Our method utilized the PanFusion model [7] to generate panoramas. However, because the PanFusion model was originally trained on simple short text prompts, it required appropriate adjustments and fine-tuning to handle complex long texts. Therefore, we added a Multi-Layer Perceptron (MLP) [9] layer and a Low-Rank Adaptation (LoRA) [10] layer to the original PanFusion model, fine-tuning it using the reannotated Habitat Matterport Dataset [11]. The MLP layer captures higher-order features in complex long texts, while the LoRA layer enhances the model’s representation ability through low-rank matrix decomposition. For model fine-tuning, we first reannotated the dataset using the generated complex long texts, pairing them with corresponding panoramas to form new training samples. Next, we froze the pre-trained parts of the PanFusion model and trained only the newly added MLP and LoRA layers. This ensures that the model effectively utilizes information from complex long texts to generate high-quality panoramas. These panoramas encapsulate the details and scenes described in the input text, serving as foundational data for subsequent steps.
Given the input long text representation $T$ and its embedding $e_{T}$, the representation after the MLP layer is:
$\mathbf{h}=\sigma(\mathbf{W}_{2}\cdot\text{ReLU}(\mathbf{W}_{1}\cdot\mathbf{e}_{T}+\mathbf{b}_{1})+\mathbf{b}_{2}),$ (2)
where $W_{1}$ and $W_{2}$ are the weight matrices of the MLP layer, $b_{1}$ and $b_{2}$ are the bias terms, and $\sigma$ is the activation function.
The representation of the LoRA layer is:
$\mathbf{h}^{\prime}=\mathbf{h}+\mathbf{U}\cdot\text{softmax}(\mathbf{V}\cdot\mathbf{h}),$ (3)
where $U$ and $V$ are the low-rank matrices of the LoRA layer.
III-B1 Dataset Reannotation
We reannotated the dataset using the generated complex long texts. Specifically, we paired the complex long texts with corresponding panoramas to form new training samples.
III-B2 Model Fine-tuning
During fine-tuning, we froze the pre-trained parts of the PanFusion model and only trained the newly added MLP and LoRA layers. This ensures that the model can effectively utilize information from the complex long texts to generate high-quality panoramas.
The adjusted and fine-tuned PanFusion model is applied to the generated long text collection, producing 360-degree panoramas. These panoramas contain details and scenes described in the input text and serve as the foundational data for subsequent steps.
III-C Panoramic Image Optimization
@@
As the pre-trained PanFusion model can only generate panoramas with a resolution of $512\times 1024$, directly using these panoramas for multi-view extraction and point cloud reconstruction would result in sparse point clouds with blurry 3D Gaussian rendering [8]. To meet the requirements of point cloud reconstruction, it is necessary to further enhance the resolution and quality of the panoramas. The specific steps are as follows: We adopted a three-step super-resolution reconstruction method involving ControlNet-Tile [12], RealESRGAN [13], and a second application of ControlNet-Tile to maximize the enhancement of image details and resolution.
III-C1 ControlNet-Tile Super-Resolution Reconstruction